OpenStudy (dls):

If the centre of mass of three particles of masses 10,20 and 30g is at the point(1,-2,3)then where should a fourth particle of mass 40g be placed,so that the combined centre of mass of the system is (1,1,1)

OpenStudy (dls):

@mayankdevnani

OpenStudy (dls):

@Yahoo!

OpenStudy (anonymous):

Position of Cm = m1x1 + m2x2 + m3x3 / (m1+m2+m3)

OpenStudy (dls):

I didn't understand the co-ordinates,is it 1,0 -2,0 & 3,0?

OpenStudy (anonymous):

There are three Co-ordinates....For the 1 St one..CM is @ (1,-2,3)

OpenStudy (dls):

10-40+90/60 =1?

OpenStudy (dls):

i didnt get it what to do

OpenStudy (anonymous):

First Find the co-ordinates of m1 m2 and m3

OpenStudy (anonymous):

(1,-2,3 ) x , y , z x axis of Cm = m1x1 + m2x2 + m3x3 / (m1+m2+m3) y axis of Cm = m1y1 + m2y2 + m3y3 / (m1+m2+m3) etc...

OpenStudy (anonymous):

u got it..)

OpenStudy (dls):

thanks :D then?

OpenStudy (dls):

@Mashy

OpenStudy (anonymous):

wait .. i know.. on call with gf :P.. 2 mins

OpenStudy (anonymous):

ok.. so let xcm be the centre of mass of the three particle system then \[Mxcm = m1x1+ m2x2 + m3x3 \] now .. when the fourth particle is added.. the centre of mass shifts hence \[M'xcm' = m1x1+ m2x2 + m3x3+ m4x4\] or \[M'xcm' = Mxcm + m4x4\] the only thing you have to find now is x4.. so calculate.. do the same in other 2 axes